OpenStudy (anonymous):

Please help!! The vertices of a triangle are A(16, 0), B(9, 2), and C(0, 0). Which side is also an altitude with a slope of -2/7? side AB side BC side AC

OpenStudy (anonymous):

AB Need explanation?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Try to draw a diagram roughly, points A and C are on the x axis. So from C to B the line will go up (positive slope), and from B to A the line will go down (negative slope). Does that makes any sense? :)

OpenStudy (anonymous):

Yupp

OpenStudy (anonymous):



OpenStudy (anonymous):

slope of AB=(2-0)/(9-16)=-2/7 AB=sqrt[(9-16)^2+(2-0)^2]=sqrt(49+4)=sqrt53 BC=sqrt[(0-9)^2+(0-2)^2]=sqrt85 AC=sqrt [(0-16)^2+(0-0)^2]=16 neither AC^2=BC^2+AB^2,Because 53+85=138 not =to 256 nor BC^2=AC^2+AB^2,Because 256+53=309 not= to 85 AB is not an altitude of triangle ABC though it has slope -2/7